OpenStudy (anonymous):

help please locate the critical points and determine the maxima and minima: 1. y=x^4+2x^3+8x+3 2. y=x^3*(x-2)^2

OpenStudy (anonymous):

differentiate and equate to zero: y' = 4x^3 + 6x^2 + 8 = 0 x = -2 is the only real solution y" = 12x^2 + 12x which is positive at x = -2 so this is a maximum turning point maximum at x=-2 y = 16 -16 -16 +3 = -13 critical point (maximum) at (-2,-13)

OpenStudy (anonymous):

you can solve 2. by the same general method

OpenStudy (anonymous):

sorry - i had a mental block !!!! a positive second derivative means a MINIMUM turning point replace maximum by MINIMUM

OpenStudy (anonymous):

i'll start number 2 for you y' = 3x^2 * (x-2)^2 + x^3*2(x-2) = 0

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

(x-2)(3x^2(x-2) + 2x^3) = 0 so one solution is x = 2 3x^3-6x^2 + 2x^3 = 0 5x^3 - 6x^2 = 0 x^2(5x - 6 ) = 0 x = 0 or x = 6/5

OpenStudy (anonymous):

so there are critical points at x=-2, x=0 and x= 6/5 now we have to find y" and plug in these values to find the nature of these points

OpenStudy (anonymous):

y' = (x-2)(5x^3- 6x^2) = 5x^4 - 6x^3 - 10x^3 + 12x^2 = 5x^4 - 16x^3 + 12x^2 y" = 20x^3 - 48x^2 + 24x at x=2 this is postive - so a minimum at x = 6/5 this is negative s0 a minimum at x = 0 this is also zero - this looks like a point of inflection MINIMUM AT (2, 2^3(2-2)^2) = (,2,0) maximum at (1.2, 1.1) point of inflection at (0,0)

OpenStudy (anonymous):

that was a long one! - i checked it out on graphical calculator to make sure its right

OpenStudy (anonymous):

ok - do you follow it alright?

OpenStudy (anonymous):

at one point i said x= -2 is one of the points - that was wrong its x = 2 - as i said later

OpenStudy (anonymous):

that second one led to some messy algebra

OpenStudy (anonymous):

thanks a bunch